062818 UCSC fog update

Although earlier this week there was more fog, today no viable sample was collected. The collector was cleaned yesterday. I took a blank sample of 160ml and then treated it with HCl.

Took more acid cleaned jars to roof

-Marcella

170831 UCSC

The fog collector came on for 3 minutes this morning at 8:43 am. No samples was collected. The collector was clean with soap and water and dried. No blank was taken.

 

(what was going on for 3 minutes on the roof?)